OpenStudy (anonymous):

Write the equation for 5x + 2y = 3 in slope-intercept form

OpenStudy (anonymous):

oh, need to find m or gradient first

OpenStudy (anonymous):

then find x

OpenStudy (anonymous):

u try first

TheSmartOne (thesmartone):

slope intercept form y=mx+b What can you do you make it into that form first?

OpenStudy (anonymous):

have u try it? y = mx + c m is also gradient or slope so try to change the position to become y = mx + c 5x+2y=3 2y=-5x+3 y=-5/2x+3/2 c=3/2 so, u get the m = -5/2 (the slope) the answer is y=-5/2x+3/2

OpenStudy (anonymous):

5X+3=2Y THEN DIVIDE TWO FROM BOTH SIDES

OpenStudy (anonymous):

i hope can help u...